OpenStudy (anonymous):

Solve for x: x^2-12x-15=0

OpenStudy (anonymous):

\[x^2-12x-15=0\] \[x^2-12x=15\] \[(x-6)^2=15+36=51\] \[x-6=\pm\sqrt{51}\] \[x=6\pm\sqrt{51}\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

m